OpenStudy (anonymous):

An automotive insurance company has 30,000 policyholders, and the accident rate is 0.06. If the company needs to make a profit of $1,740,000 this year and the payout for each claim (after deductibles) is $8,200, the minimum premium it must charge is $_______. The company’s financial advisers have predicted that in the upcoming year, the accident rate will increase by 5% of the base rate and the number of policyholders will increase by 12%. If the company’s profit needs and the payout for each claim remain the same, the minimum premium should increase by $_____.

OpenStudy (anonymous):

@Preetha @mathmate

OpenStudy (welshfella):

First part:- Profit = income - outgoings outgoings = 30,000 * 0.06 * 8200 dollars profit = 1,740,000 income = 30,000 x where x = premium

OpenStudy (welshfella):

form an equation and solve for x

OpenStudy (anonymous):

550?

OpenStudy (welshfella):

yes good job

OpenStudy (anonymous):

thanks!

OpenStudy (welshfella):

yw for the second part you need to multiply 0.06 by 1.05 ( = 5*% increase)

OpenStudy (welshfella):

and the 30,000 by 1.12 ( = 12% increase) and do the same calculation

OpenStudy (anonymous):

wait would you add them?

OpenStudy (welshfella):

no the calculation with the 2 above changes will give a new premium they want the increase in premium so subtract the iold premium from the new one

OpenStudy (anonymous):

33599.937?

OpenStudy (welshfella):

what is that figure?

OpenStudy (anonymous):

I subtracted the (.06*1.05) from the (30,000*1.12)

OpenStudy (anonymous):

did i do it wrong?:(

OpenStudy (welshfella):

this is way to work out the new premium outgoings = 30,000 * 1.12* 0.06 * 1.05 * 8200 dollars profit = 1740000 income = 30000*1.12* x x = (174000 + 30,000 * 1.12* 0.06 * 1.05 * 8200 ) / (30,000 * 1.12)

OpenStudy (welshfella):

x = 568.39 dollars so increase = 568.39 - 550

OpenStudy (welshfella):

yes that subtraction was wrong

OpenStudy (anonymous):

so 18.39 is the answer?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

ok thank you!

OpenStudy (welshfella):

you have to go back to the original formula with the new amounts

OpenStudy (welshfella):

yw